Item 9 Item 9 Schager Company purchased a computer system on January 1, 2019, at a...

Item 9

Item 9

Schager Company purchased a computer system on January 1, 2019, at a cash cost of $34,000. The estimated useful life is 10 years, and the estimated residual value is $3,900. The company will use the double declining-balance depreciation method.

What is the accumulated depreciation balance as of December 31, 2020?
Answer #1

Double declining rate = 200/useful life = 200/10 = 20%

Depreciation for year 2019 = 34,000*20% = 6800

Depreciation for year 2020 = (34,000-6800)*20% = 5440

Accumulated Depreciation as of Dec 31,2020

= 6800+5440

= 12240